# VP-MEL: Visual Prompts Guided Multimodal Entity Linking

### v2

@@ 5.3 Detailed Analysis
As shown in Table 3, without mention words, the performance decreases significantly across three metrics. The average proportion of performance drop across three metrics is 72.65%, 64.48%, and 60.28%, respectively. The inability to extract meaningful information from both visual and textual data shows that MEL methods are ill-suited for tasks that lack mention words. We use FBMEL from the VP-MEL task to test on WikiDiverse. After removing the Detective-VLM, and in the absence of both visual prompts and mention words, FBMEL can still achieve metrics of 23.87%, 38.37%, and 45.14%. This demonstrates that FBMEL in the VP-MEL has a stronger ability to utilize information from both image and text.
@@
As shown in Table 4, to analyze the effectiveness, we replace the Detective-VLM with various VLMs for experiments and obtain the results. Our method achieves the best metrics. Compared with the second best metric, Detective-VLM gains 5.89% absolute improvement of Hit@1. The output of Detective-VLM is concise and effective, allowing it to perform better within FBMEL. Non-fine-tuned VLMs usually generate a large amount of irrelevant information, which interferes with subsequent usage.
@@
Table 5: Comparison of performance across different feature layers in $V^{G}$.
@@
In Table 6, we conduct ablation study of FBMEL framework. First, the Detective-VLM is removed from FBMEL. All metrics exhibit a downward trend. Even without additional textual information, FBMEL can still effectively perform entity linking. Moreover, without the Detective-VLM, FBMEL still outperforms MEL methods. This indicates that FBMEL is capable of leveraging information from both images and text more efficiently than MEL methods. To further evaluate the contribution of Visual Prompts, we removed the visual prompts from the images based on the previous setup. All metrics show decreases. This indicates that visual prompts help the model focus on specific regions in images. The function of visual prompts is significantly enhanced by the Detective-VLM, which facilitates the ability to derive more comprehensive information from images.
@@ 6 Conclusion
In this paper, we introduce a new task VP-MEL and construct a new dataset VPWiki, aiming to utilize visual prompts to broaden application scenarios of MEL. To address the VP-MEL task, we propose FBMEL, a method designed to effectively leverage visual prompts for extracting meaningful information from image-text data. Extensive experimental results demonstrate that FBMEL surpasses state-of-the-art methods, effectively reducing dependency on textual information while significantly improving the utilization of visual information. VP-MEL significantly reduces the constraints of mention words and expands the applicability of MEL to real-world scenarios.
@@ A.3 Evaluation Metrics
$\mathrm{Accuracy}_{top-k}=\frac{1}{N}{\sum_{i}^{N}I(t_{i}\in y_{i}^{k})}$
@@ A.4 Detective-VLM
Detective-VLM is based on mplug-owl2 (Ye et al., 2024), and we perform instruction fine-tuning on mplug-owl2-llama2-7b  [^9] https://huggingface.co/MAGAer13/mplug-owl2-llama2-7b.
We utilize VPWiki dataset to design the fine-tuning dataset, where $\{Image\}$ and $\{Sentence\}$ correspond to $m_{v_{j}}$ and $m_{s_{j}}$ in $M_{j}$, respectively. In the fine-tuning dataset, the ${Entity\ \allowbreak Name}$ corresponds to the mention words in $M_{j}$ that are associated with the Visual prompt, the $\{Entity\ \allowbreak Type\}$ is one of [$Person$, $Organization$, $Location$, $Country$, $Event$, $Works$, $Misc$].
@@ A.6 Pipeline Module Settings
The construction of the pipeline based on the Visual Entailment Module and Visual Grounding Module is inspired by Li et al. (2024). In the pipeline, Visual Entailment Module is employed to evaluate the relationship between the mention words and the corresponding images. After filtering out the highly relevant data, Visual Grounding Module annotates the grounding box in the images. For the Visual Entailment Module and Visual Grounding Module, we choose $\mathrm{OFA}_{large(VE)}$ and $\mathrm{OFA}_{large(VG)}$ (Wang et al., 2022b), respectively.
@@ A.7 WikiDiverse and WikiMEL
During the data collection process, we selected all of the WIKIDiverse data and 5,000 samples from the WikiMEL dataset. Compared to WikiMEL, the images in the WIKIDiverse dataset are more content-rich and more representative of real-world application scenarios, making them better suited to the needs of the VP-MEL task in practical applications. Therefore, WIKIDiverse constitutes a larger portion of the VPWiki dataset. We integrate the KB from both datasets. The entity set consists of all the entities in the main namespace with the size of 28M.
@@ A.8 Additional Experiments and Experimental Descriptions
The knowledge base (KB) used in the WikiMEL and RichpediaMEL datasets differs from that of VPWiki and WIKIDiverse. To thoroughly evaluate the performance of the FBMEL framework in the MEL task, we retrained FBMEL on the WikiMEL and RichpediaMEL datasets. Notably, during the retraining process, neither visual prompts nor the Detective-VLM are used, meaning that in this experiment, FBMEL can only be regarded as a MEL model. The experimental results are shown in Table 8. FBMEL is second only to MIMIC and MELOV, demonstrating its competitive performance as an MEL method.
(a) Correct cases.
(b) Error cases.
@@ A.9 Case Study
To clearly demonstrate the proposed VP-MEL task and the FBMEL model, we present case studies and compare them against two strong competitors (i.e., LLaVA-1.5 and MIMIC), in Figure 6. As shown in Figure 6(a), in the first case, all three methods correctly predicted the entity. FBMEL makes full use of both image and text information, allowing it to more effectively distinguish between the different individuals in the image. LLaVA-1.5 may be overwhelmed by the textual information, while MIMIC struggles to identify the correct entity when the mention words are unavailable. In the second case, both LLaVA-1.5 and MIMIC retrieve Endeavour as the first choice. Only FBMEL, with the guidance of Visual Prompts and integration of textual information, correctly identified the right entity. In Figure 6(b), we present the error cases. In the first case, when the content of the image interferes with the visual prompt, it impairs FBMEL’s reasoning process. The red box in the image bears a high similarity to the visual prompt. As a result, FBMEL incorrectly focuses on the wrong region of the image, ranking Donald Trump first. When FBMEL encounters difficulties in distinguishing the objects within the image’s visual prompts, it leads to incorrect inferences. For example, in the second case, the distinguishing features of the two individuals in the image are obstructed, causing FBMEL to encounter difficulties in differentiating between the two. The image content in real-world data is often complex, making VP-MEL a task that presents significant challenges. We hope that this task can be further refined and developed over time.

### v3

@@ 6.3 Detailed Analysis
For a fair comparison, we remove the Detective-VLM module from FBMEL. As shown in Table 3, the performance of MEL methods drop significantly across all three metrics in the absence of mention words. The average performance decline is 72.65%, 64.48%, and 60.28%, respectively. This indicates that MEL methods fail to extract meaningful information from visual and textual data, making them unsuitable for tasks without mention words. In contrast, even without Detective-VLM, visual prompts, or mention words, FBMEL can still achieve the best metrics. This demonstrates that FBMEL in the VP-MEL task possesses a stronger capability to leverage both image and text information effectively.
@@
As shown in Table 4, we evaluate the effectiveness of Detective-VLM by replacing it with various VLMs and analyzing the results. Our method achieves the best performance across all metrics. In particular, Detective-VLM shows an absolute improvement of 5.89% in Hit@1 compared to the second-best approach. In contrast, non-fine-tuned VLMs often produce a large amount of irrelevant information, which hampers subsequent processing.
@@
Table 5: Performance comparison across different feature layers in $V^{G}$.
@@
In Table 6, we conduct ablation study on the FBMEL framework. First, we remove the Detective-VLM module from FBMEL, which results in a decline across all metrics. Notably, even without Detective-VLM, FBMEL shows robust entity linking performance, outperforming MEL methods as shown in Table 2(a). This highlights the ability of FBMEL to efficiently leverage multimodal information from both images and text. To further evaluate the impact of Visual Prompts, we remove the visual prompts from the images in the previous setup. This leads to a noticeable drop in all metrics, underscoring the importance of visual prompts in guiding the model to focus on relevant regions within images.
@@ 7 Conclusion
In this paper, we introduce a new task, VP-MEL, aiming to link visual regions in image-text pairs to the corresponding entities in knowledge base, guided by visual prompts. To support this task, we construct a high-quality dataset, VPWiki, leveraging an automated annotation pipeline to enhance the efficiency of data annotation. To address the VP-MEL task, we propose FBMEL, a framework that can effectively utilize visual prompts to extract enhanced local visual features and generate supplementary textual information. FBMEL maintains a balance between visual and textual features, preventing excessive reliance on a single modality. Extensive experimental results demonstrate that FBMEL surpasses state-of-the-art methods. Moreover, VP-MEL significantly alleviates the constraints of mention words and expands the applicability of MEL to real-world scenarios.
@@ A.3 Evaluation Metrics
$\mathrm{Accuracy}_{top-k}=\frac{1}{N}{\sum_{i}^{N}I(t_{i}\in y_{i}^{k})},$
@@ A.4 Detective-VLM
Detective-VLM is based on the mplug-owl2 framework (Ye et al., 2024), with instruction fine-tuning carried out using the mplug-owl2-llama2-7b model  [^9] https://huggingface.co/MAGAer13/mplug-owl2-llama2-7b.
We utilize VPWiki dataset to design the fine-tuning dataset, where $\{Image\}$ and $\{Sentence\}$ correspond to $m_{v_{j}}$ and $m_{s_{j}}$ in $M_{j}$, respectively. In the fine-tuning dataset, the $\{Entity\ \allowbreak Name\}$ corresponds to the mention words in $M_{j}$ that are associated with the Visual prompt, the $\{Entity\ \allowbreak Type\}$ is one of [$Person$, $Organization$, $Location$, $Country$, $Event$, $Works$, $Misc$].
@@ A.6 Pipeline Module Settings
For the Visual Entailment Module and Visual Grounding Module, we choose $\mathrm{OFA}_{large(VE)}$ and $\mathrm{OFA}_{large(VG)}$ (Wang et al., 2022b), respectively.
@@ A.7 WikiDiverse and WikiMEL
During the data collection process, we select the entire WIKIDiverse dataset along with 5,000 samples from the WikiMEL dataset. Compared to WikiMEL, WIKIDiverse features more content-rich images that better represent real-world application scenarios, making it particularly suitable for meeting the requirements of the VP-MEL task in practical contexts. Consequently, WIKIDiverse constitutes the majority of the VPWiki dataset. Additionally, we integrate the knowledge bases (KBs) from both datasets, resulting in an entity set encompassing all entities in the main namespace.
@@ A.8 Additional Experiments and Experimental Descriptions
The knowledge base (KB) used in the WikiMEL and RichpediaMEL datasets differs from that of VPWiki and WIKIDiverse. To comprehensively assess the performance of the FBMEL framework in the MEL task, we retrain FBMEL on the WikiMEL and RichpediaMEL datasets. Importantly, neither visual prompts nor the Detective-VLM are employed during the retraining process. The experimental results are shown in Table 8. FBMEL is second only to MIMIC and MELOV, demonstrating its competitive performance as an MEL method.
(a) Successful predictions.
(b) Failed predictions.
@@ A.9 Case Study
To clearly demonstrate the proposed VP-MEL task and the FBMEL model, we conduct case studies and compare them against two strong competitors (i.e., LLaVA-1.5 and MIMIC), in Figure 6. As shown in Figure 6(a), in the first case, all three methods correctly predicted the entity. FBMEL makes full use of both image and text information, allowing it to more effectively distinguish between the different individuals in the image. LLaVA-1.5 may be overwhelmed by the textual information, while MIMIC struggles to identify the correct entity when the mention words are unavailable. In the second case, both LLaVA-1.5 and MIMIC retrieve Endeavour as the first choice. Only FBMEL, with the guidance of Visual Prompts and integration of textual information, correctly predicts the right entity. In Figure 6(b), we present the failed predictions. In the first case, when the content of the image interferes with the visual prompt, it impairs the reasoning process of FBMEL. The red box in the image bears a high similarity to the visual prompt. As a result, FBMEL incorrectly focuses on the wrong region of the image, ranking Donald Trump first. When FBMEL encounters difficulties in distinguishing the objects within the visual prompts, it leads to incorrect inferences. For example, in the second case, the distinguishing features of the two individuals in the image are obstructed, which causes FBMEL to struggle in differentiating between them. The image content in real-world data is often complex, which makes VP-MEL a challenging task. We hope that this task can be further refined and developed over time.